### v1

@@ Abstract
Exercising regularly is widely recognized as a cornerstone of health, yet the challenge of sustaining consistent exercise habits persists. Understanding the factors that influence the formation of these habits is crucial for developing effective interventions. This study utilizes data from Mars Athletic Club, Türkiye’s largest sports chain, to investigate the dynamics of gym attendance and habit formation. The general problem addressed by this study is identifying the critical periods and factors that contribute to the successful establishment of consistent exercise routines among gym-goers. Here we show that there are specific periods during which gym attendance is most crucial for habit formation. By developing a survival metric based on gym attendance patterns, we pinpoint these critical periods and segment members into distinct clusters based on their visit patterns. Our analysis reveals significant differences in how various subgroups respond to interventions, such as group classes, personal trainer sessions, and visiting different clubs. Using causal inference analysis, we demonstrate that personalized guidance and social dynamics are key drivers of sustained long-term engagement. By systematically examining these variables and considering the specific characteristics of different clusters, our research demonstrates the importance of a tailored, multi-dimensional approach to promoting exercise habits, which integrates social dynamics, personalized guidance, and strategic interventions to sustain long-term engagement.
@@ 1 Introduction
Many people struggle to make exercise a consistent part of their lives, despite the well-documented benefits of physical activity. Regular exercise is associated with numerous health advantages, including improved cardiovascular health, enhanced mental well-being, and a reduced risk of chronic diseases such as obesity, diabetes, and certain cancers [1, 2]. The World Health Organization estimates that nearly 25% of adults globally do not meet the recommended levels of physical activity, which contributes significantly to the rising prevalence of chronic conditions like obesity, cardiovascular disease, and diabetes [3]. Moreover, approximately 9% of premature deaths are attributed towards sedentary lifestyles globally, which shows the critical importance of encouraging more widespread participation in regular exercise [4].
This gap between awareness of the benefits of exercise and the actual adoption of a regular workout routine is influenced by a complex interplay of behavioral, psychological, demographic, and environmental factors [5, 6, 7]. Psychological barriers, such as present bias and time inconsistency often lead individuals to favor immediate comforts over long-term health benefits, resulting in procrastination and a lack of consistent exercise [8]. Additionally, the process of transforming sporadic exercise into a consistent routine generally unfolds in two phases: the initiation phase, driven by immediate motivations like health concerns or aesthetic goals, and the execution phase, where exercise becomes ingrained in one’s lifestyle [9]. In addition, consistent routines, simple exercise activities, and positive emotional experiences are significant predictors of successful habit formation. [10]. Milkman also talks about the effect of positive emotional experiences on habit formation, who suggests that temptation bundling is a useful strategy that combines the tasks one needs to do with those that one enjoys. For instance, listening to a favorite podcast only while working out can help overcome the inertia of undesirable habits and foster long-term adherence to beneficial routines like regular exercise [11].
Small incentives can play a surprisingly large role in encouraging exercise adherence. For instance, research has shown that financial incentives can increase gym attendance, especially for those who did not previously attend regularly [12]. One study found that rewarding participants for coming back to the gym after they had missed a previous workout paved the way for an increase of 0.4 more weekly gym visits, with just a bonus of 125 points that would be worth almost 10 cents [13]. These findings align with broader research indicating that even very small incentives can have a disproportionately large impact on behavior [14].
Moreover, social dynamics play a critical role in reinforcing exercise habits. Group exercises, for example, provide motivational feedback and foster a sense of community, which can drive individuals to increase their workout intensity and frequency [15, 16, 17]. Small-group, individualized personal training sessions can support basic psychological needs, autonomous exercise motivation, and exercise self-efficacy [18]. Peer influence and social norms are also powerful factors that can significantly impact how regularly and intensely people exercise, which demonstrates that people’s behaviors are often shaped by the actions and expectations of those around them [19, 9].
Personalizing exercise programs is also a crucial factor in encouraging long-term physical activity. Studies have shown that personalized training sessions with certified personal trainers help maintain exercise behavior and motivation, particularly in groups such as female college students [20]. Furthermore, supervised, periodic exercise programs, which are often customized to the individual’s fitness level and goals, have been shown to lead to greater improvements in physical fitness and adherence compared to self-directed training [21]. This is because the presence of a trainer provides both accountability and expert guidance, which are critical for sustaining motivation over time. Additionally, supportive interactions from exercise instructors, who can offer personalized feedback and encouragement, have been found to promote long-term exercise participation in gym settings significantly, further emphasizing the importance of personalization in fitness programs [22]. Personalized exercise programs that combine moderate and high-intensity training are more effective at improving cardiorespiratory fitness and metabolic health than standardized programs [23]. This suggests that tailoring the intensity and type of exercise to an individual’s needs can yield better health outcomes than a one-size-fits-all approach. Moreover, resistance training with a personal trainer often results in participants selecting higher exercise intensities compared to when they train alone, demonstrating the motivational impact of personalized supervision [24]. The ability of personal trainers to adjust workout intensity based on real-time feedback helps individuals push their limits safely, thereby achieving better results.
Despite the comprehensive understanding of these individual factors, there remains a notable gap in the literature concerning holistic intervention models that integrate these insights into cohesive, computational frameworks for practical application. Current research often focuses on isolated aspects of behavior modification without a unified model that accounts for the complex interplay of psychological, social, and environmental influences in a real-world setting. Our research utilizes comprehensive data from Mars Athletic Club, Türkiye’s largest sports chain. First, we developed a survival metric based on the frequency and consistency of gym visits to define what constitutes a lasting exercise habit. This metric is a foundational tool to assess the transformation from sporadic to regular exercise patterns among gym-goers. Additionally, we identified distinct clusters of gym members based on their visit patterns, which revealed significant differences in how various subgroups respond to interventions. With this survival metric and cluster-based segmentation established, we applied causal inference methodologies to precisely analyze the impact of various interventions. Propensity score matching allows us to control for confounding variables and isolate the effects of factors such as participation in group classes, personal trainer sessions, and social interactions like inviting a friend to the gym. By systematically examining these variables and considering the specific characteristics of different clusters, our research aims to discover which strategies are most effective in enabling consistent exercise behaviors. The ultimate goal is to integrate these findings into actionable strategies that can be implemented by gyms and health clubs to enhance public health outcomes.

### v2

@@ Abstract
Regular exercise is widely recognized as a cornerstone of health, yet sustaining consistent exercise habits remains challenging. Understanding the factors that influence the formation of these habits is crucial for developing effective interventions. This study utilizes data from Mars Athletic Club, Türkiye’s largest sports chain, to investigate the dynamics of gym attendance and habit formation. The general problem addressed by this study is identifying the critical periods and factors that contribute to the successful establishment of consistent exercise routines among gym-goers. We show that specific periods of attendance are most crucial for habit formation. By developing a survival metric based on gym attendance patterns, we pinpoint these key phases and segment members into distinct clusters based on their visit patterns. Our analysis reveals significant differences in how various subgroups respond to interventions, such as group classes, personal trainer sessions, and visiting different clubs. Using causal inference analysis, we demonstrate that personalized guidance and social dynamics are key drivers of sustained long-term engagement. By systematically examining these variables and considering the specific characteristics of different clusters, our research highlights the importance of a tailored, multi-dimensional approach to promoting exercise habits, which integrates social dynamics, personalized guidance, and strategic interventions to sustain long-term engagement.
keywords:
Habit formation , Causal inference , Non-negative Matrix Factorization , Exercise , Public health , Behavioral clusters
journal: Expert Systems with Applications
[…]
{highlights}
We define a novel survival metric to measure the effect of gym attendance on habit formation.
We identify critical 6-week and 17-week milestones for forming exercise habits.
Gym members are clustered into 5 distinct behavioral groups using NMF.
Personalized training and social dynamics are key drivers of long-term engagement.
Tailored, multi-dimensional interventions are crucial for different member groups.
@@ 1 Introduction
Developing consistent exercise habits is one of today’s most pressing public health challenges, contributing to widespread inactivity and preventable chronic disease. Regular physical activity confers well‐established health benefits—improved cardiovascular and metabolic functioning, enhanced mental well‐being, and reduced risk of noncommunicable diseases—yet fewer than 25% of adults worldwide meet recommended activity levels, and nearly 9% of premature deaths are linked to sedentary lifestyles (World Health Organization, 2018; Lee et al., 2012). Despite widespread awareness of these benefits, many people struggle to translate their intentions into enduring exercise habits. One key reason is that starting an exercise routine requires consistent motivation, while maintaining that routine relies more on automatic processes. With repeated performance in stable environments, situational cues can trigger the behavior with little conscious effort. This phenomenon is known as cue-triggered automaticity (Wood and Rünger, 2016; Gardner et al., 2022).
Early work on habit formation showed that simple health behaviors (e.g. drinking water) reach plateaued automaticity in around 66 days of daily repetition, but more complex activities (like a full workout) take longer and vary widely across individuals (Lally et al., 2010). In the exercise domain, longitudinal tracking of new gym members revealed that attending roughly four sessions per week for at least six weeks is the minimum “dosage” needed to establish a self‐sustaining gym habit—after which attendance becomes less dependent on conscious planning and more driven by contextual triggers (Kaushal and Rhodes, 2015). However, mere repetition alone is not enough: positive affect experienced during workouts significantly boosts habit strength, whereas monotonous or aversive sessions fail to produce lasting automaticity (Weyland et al., 2020). Conversely, individuals who abandon fitness trackers often cite lower internalized exercise motives and guilt about unmet goals, underscoring the need to cultivate intrinsic rather than purely extrinsic drivers during the habit-building phase (Nuss and Li, 2021). In the same vein, fitness-app users who report stronger “exercise identity” are substantially more active, suggesting that digital tools foster persistence by reinforcing self-concept as an exerciser (Barkley et al., 2020). Recent machine‐learning analyses of millions of gym visits confirm that as exercise is consistently repeated in the same environment, individuals become less context‐sensitive—that is, their behavior shifts from deliberative goal pursuit toward habit‐driven action (Buyalskaya et al., 2023). Recent evidence from wearable-technology users also shows that gamified tracking, instructional feedback, and progress sharing can accelerate the transition from intention to routine by strengthening self-determined motivation for physical activity (Oc and Plangger, 2022)
Translating these insights into real‐world interventions requires moving beyond isolated strategies to comprehensive, multi‐level models. Socioecological frameworks posit that behavior is shaped simultaneously by individual traits, social networks, physical environments, organizational settings, and policy contexts (Sallis et al., 2006). In practice, effective programs combine individual‐level techniques (e.g. skill training, self‐monitoring) with supportive social structures (group classes, workout partners), environmental cues (gym access, prompts), and even institutional policies (workplace wellness incentives) to create conditions under which exercise habits can form and endure (Sallis, 2018). Moreover, talking about workouts both one-to-one with close contacts and in wider network broadcasts exerts distinct, additive influences on physical-activity levels, underscoring the need for interventions that enable both interpersonal and group communication (Li and Meng, 2023). In online fitness communities, simply having a larger core network of active peers—and comparing one’s progress with fitter friends—predicts higher physical-activity levels, illustrating the motivational leverage of social ties (Huang et al., 2022). The COM‐B system further organizes this complexity by identifying three core determinants of behavior (capability, opportunity, and motivation) and mapping them onto intervention functions such as education, environmental restructuring, and enablement (Michie et al., 2011). Integrated applications of socioecological and COM‐B models, for instance in office‐based physical activity initiatives, demonstrate how modifying workspace design, social norms, and individual skill can jointly enhance both the contextual cues and motivational drivers necessary for habit formation (van Kasteren et al., 2020).
Despite these theoretical advances, empirical studies rarely combine habit‐formation principles, multi‐level intervention design, and rigorous longitudinal analysis in authentic gym contexts. Most research focuses either on short‐term habit psychology under controlled conditions or on community‐level interventions without explicitly measuring automaticity. Meanwhile, survival‐analysis techniques offer powerful tools to model “time‐to‐habit” and persistence: by treating the establishment of a consistent workout pattern as a time‐to‐event outcome, one can quantify the hazard of relapse (dropout) and identify when habits are most vulnerable (Harris and Kessler, 2019; Galvim et al., 2019). A recent meta‐analysis of health‐behavior habit formation confirms that while 2–3 months of consistent action is typically required to reach plateaued habit strength, individual trajectories vary greatly and are influenced by factors such as routine timing, enjoyment, and environmental cues (Singh et al., 2024).
Unlike laboratory habit studies that track reaction times or self‐report automaticity scales (Gardner et al., 2022, 2012), our large‐scale gym data lack those direct habit measures, so we use a survival‐based proxy for persistence instead. To our knowledge, no real‐world study to date has (a) segmented exercisers by temporal habit patterns, (b) mapped comprehensive, multi-level interventions onto those patterns, and (c) used time‐to‐habit survival models to quantify persistence. Habit strength is, at heart, persistence despite fluctuating motivation. Survival analysis models this directly by linking dropout risk to early cue–behavior repetition, turning habit formation into a time-to-event problem.
In this study, we bridge these gaps by using a large, real‐world dataset from Türkiye’s largest gym chain. We (1) define and apply a survival‐analysis–based metric to capture the persistence of gym attendance over time, (2) segment members into behavioral clusters based on their temporal visit patterns, and (3) employ causal‐inference methods to estimate the effects of diverse interventions—group classes, personal training, facility variety, and social invitations—while adjusting for demographic and contextual confounders. By integrating habit‐formation theory, multi-level intervention frameworks, and survival modeling, our work illuminates how and when exercise habits crystallize in a naturalistic fitness setting, and offers practical guidance for designing layered interventions that transform occasional gym‐goers into steady exercisers.
2 Methods
2.1 Dataset
Throughout this project, we analysed anonymized data provided by Mars Athletic Club. Mars Athletic Club (MAC) is the biggest gym chain in Türkiye, and has over 100 clubs in 13 provinces in the country, with most being in Istanbul, the most populated province. Founded in 2007, the number of locations has been growing steadily over the years.
Due to the COVID-19 pandemic, many countries implemented drastic measures. Türkiye was one of the countries that implemented lockdowns in 2020 and 2021, meaning that people experienced lockdowns and were unable to exercise indoors. These measures affected a lot of businesses, one of which was the fitness industry. As people were not allowed outside for extended durations, their gym memberships became obsolete during the pandemic.
Given the exceptional circumstances surrounding the data, we restricted our analysis to data from 2022 through the end of 2023. Following the insights of project partners at MAC, we concluded that it would be better if we considered only the first paid contract for all customers. This was done to account for users who may have been on subsequent contracts during this time frame, and this would mean that they would already have had a chance to achieve some habit formation before, which would give them an advantage over other members in forming habits. Additionally, we agreed to limit the contract type to annual, instead of having different contract types such as 6-month or monthly contracts. This would give every user an equal opportunity to either form habits within a reasonable time frame or fail to do so. Furthermore, we also filtered out non-paid contracts because these were usually allocated to employees. Crucially, all data used in this study were fully anonymized to safeguard personal information in compliance with Türkiye’s Personal Data Protection Law (KVKK) (Personal Data Protection Authority, 2016). This anonymization ensures the privacy of MAC’s members while allowing for a comprehensive analysis of gym usage trends and behaviors without compromising individual confidentiality. No identifiable information was available to the research team at any stage of the analysis.
The variables used to analyse people’s habitual differences are listed below.
  • 1.
    Main club: The club that the contract was signed at.
  • 2.
    Membership Category: Membership plan associated with the customer, which varies by club type and package.
  • 3.
  • 4.
    Number of sessions with a personal trainer: The number of times that a customer worked with a personal trainer. These workouts are sold as sets of 10, and the customer may choose to train with any trainer they wish.
  • 5.
  • 6.
    Number of different group classes attended: The number of distinct group class categories (e.g., GFX, core, cardio) attended by a customer.
  • 7.
    Number of invitation credits used: Total number of credits that the customer has used to invite their friends for a single gym session. Different numbers of credits are issued for different membership categories every month.
  • 8.
    Previous experience level: Self-reported prior experience level of the customer with fitness from MAC+ mobile application. Levels of 0 (beginner) through 3 (experienced). Available for around 40.2% of the customers.
  • 9.
    Form Level: Self-reported current form level of the customer from MAC+ mobile application. For males, it is measured as the number of push-ups they can do in a minute, and for females, it is the number of bodyweight squats they can do in a minute. Levels of 0 (unfit) through 2 (very fit). Available for approximately 33.6% of the customers.
  • 10.
    Estimated Frequency of Exercise: Self-reported estimated frequency of exercise from MAC+ mobile application. Levels of 0 (no exercise) through 2 (regular exercise). Available for more than 40% of the customers.
  • 11.
2.2 Clustering Customers
We hypothesize that users may differ in their exercise habits, and these differences are often attributed to age and gender. However, external factors like exercise environment and daily routines also determine when people find opportunities for these activities. We used Non-negative Matrix Factorization (NMF) to identify distinct behavioral groups when studying daily and weekly gym visits. Different groups may have varying tendencies and independent habit formations, which we could incorporate to our analyses. The motivation for using NMF is its inherent ability to cluster data and its ease of interpretation compared to other clustering methods (Ding et al., 2005). Furthermore, with this method, one can easily infer the cluster membership of input data. The NMF algorithm approximates the factorization of the data matrix $\mathbf{V}$ onto two non-negative matrices, $\mathbf{W}$ and $\mathbf{H}$, as shown in Equation 1.
$\mathbf{V}\approx\mathbf{W}\mathbf{H}$ (1)
where $\mathbf{V}\in\mathbb{R}_{+}^{m\times n}$ is the matrix with rows corresponding to features and columns to customers (observations), $\mathbf{W}\in\mathbb{R}_{+}^{m\times k}$ is the feature (basis) matrix, and $\mathbf{H}\in\mathbb{R}_{+}^{k\times n}$ is the coefficient matrix (Kim and Park, 2008) with $k$ as a pre-specified parameter, denoting the number of components to represent the data matrix. Using the coefficient matrix $\mathbf{H}$, we assign data point $v_{j}$ to cluster $k$ if $\mathbf{H}_{kj}$ is the maximum across all components. To cluster the customers using NMF, we vectorized all the visits of each customer during their first six weeks by grouping them by the day of week and hour of day. This transforms each customer’s visit history into a vector where each element corresponds to a specific day-of-week and hour-of-day time slot. The vectorization process is explained in detail in Appendix B.
As the first step, we conducted customer clustering on a single club that was representative of the whole customer base. Accordingly, we analysed one of the most popular MACFit branches in Istanbul, which had around 4,500 members at the time of the analysis. The company suggested this club, as they knew it resembled the general customer population. We experimented with a few values for $K\in\{3,\dots,8\}$ to select the optimal number of clusters, and after $K=5$, we observed diminishing returns in terms of reconstruction error. Additionally, larger values of $K$ produced clusters that were less interpretable. We thus settled on 5 clusters. Because the five-cluster solution separated the visiting patterns well, we applied NMF to the full customer base. Our results showed that 5 clusters were consistent throughout our whole customer base, thereby confirming the choice of $K$. The visit patterns of the customers are presented in Fig2(a).
Initially, customers were categorised strictly, with each customer assigned to exactly one cluster. However, later on, since we would be using propensity score matching (PSM), having the probabilities of belonging to each cluster would provide significant benefits for the matching algorithm. Thus, we utilized the output of the NMF algorithm to get the probabilities. The coefficient matrix $\mathbf{H}$ can be used either for hard assignment of customers to clusters or for calculating membership probabilities. We used a softmax function across the $K$ components for each customer to calculate the probability of belonging to the clusters as follows.
$P(z_{j}=k)=\frac{\exp(\mathbf{H}_{kj})}{\sum_{i=1}^{K}\exp(\mathbf{H}_{ij})}$
This normalization ensures a valid probability distribution, with values between 0 and 1 that sum to 1 across all $K$ clusters for each customer. The resulting vectors capture how strongly each of the $K$ visit patterns contribute to each customer’s visit profile. The cumulative distribution functions for the probabilities of weeks 6 and 17 are presented in Figures 2(c,d).
2.3 Cluster-level Deviations for Demographic Groups
In each behavioral cluster, user demographics may differ, with certain age and gender groups more likely to be observed in some clusters. To distinguish the prior probability of observing certain user demographics ($D_{i}$) from their preferential occurrences, we calculated the likelihood ratios between conditional probability of being in a given cluster ($C_{i}$) following Eq. 2
$\text{Likelihood ratio for demographic $D_{i}$ in cluster $C_{j}$}=\frac{P(D_{i}\mid C_{j})}{P(D_{i})}$ (2)
This value will be 1 if cluster assignment has no effect on observing certain demographic groups. A deviation from 1 indicates whether a group is more (>1>1) or less (<1<1) likely to be observed. We obtained the deviations by subtracting 1 from this value.
2.4 Defining a Survival Metric for Habit Formation
Understanding the underlying mechanisms of habit formation is key to developing targeted interventions that motivate individuals to attend the gym consistently. We developed the survival metric to measure and analyze the persistence of gym habits. Our approach to this analysis was inspired by the work of Harris and Kessler (2019), who explored habit formation and activity persistence in the context of gym equipment usage. Their study demonstrated that frequent early activity leads to more persistent exercise behavior, suggesting that interventions aiming for behavioral change need to involve higher frequency or longer duration activities to be effective (Harris and Kessler, 2019). This metric was defined as the number of consecutive weeks in which a member attended the gym at least once. We added a tolerance of one week of absence called gap week, acknowledging that occasional interruptions (e.g., travel or illness) should not be taken as a disruption of habit formation. The decision to include this tolerance was based on an analysis of intermediate periods between consecutive gym attendance periods, which can be found in Appendix C, which showed that a one-week gap was most commonly associated with sustained attendance.
However, due to the inclusion of this tolerance period, the number of weeks it took for individuals to reach each survival milestone (i.e., 6, 7, …,17 weeks) varied across users. As a result, when creating attendance milestones, it was imperative to assess users’ habit formation based on the number of visits made during a fixed period rather than the survival longevity. This approach also provided a fixed period for later analysis of users’ behaviors and allowed us to investigate the impact of early actions on long-term persistence.
To further assess and track habit formation, we established specific attendance milestones corresponding to critical survival durations. These milestones mark a member’s progress in their gym attendance journey. Each milestone is defined by a critical visit count that members must achieve within the specified period. For example, to reach the $\text{6}^{\text{th}}$ week milestone, a member must attend the gym at least 9 times within the first six weeks of their membership. Figure 4(d) shows an almost linear trend in the critical number of visits required to reach milestones. We fitted a linear model y=2.01x5.35y=2.01x-5.35, with $\text{slope}\approx 2.01$, indicating that members need to visit the gym roughly twice per week to meet their milestones. This suggests that attending the gym at least twice a week is important for maintaining consistent habits. These milestones were introduced to provide a structured approach for evaluating habit formation over time. Users may still attend the gym for extended periods, but these milestones allow us to separate users with the same length of membership into two categories. By focusing on these milestones, we aimed to simplify the tracking of member progress and to identify points at which interventions might be most effective.
To determine the critical minimum number of visits required to reach a 6-week survival milestone, we compared the distribution of the number of visits made in 6 weeks by the members who survived 6 weeks or less and who survived more than 6 weeks in Figure 1(a). We analyzed the cumulative distribution of visits for both groups, visualized in Figure 1(b), examined the difference between the cumulative distribution functions (CDFs), and determined the number of visits that maximized the difference between the two groups, as shown in Figure 1(c). The critical number of visits was nine within the first six weeks, as this threshold maximized the difference between those who maintained their gym habits and those who did not. This approach allowed us to identify the number of visits that best demonstrated sustained engagement and habit formation during the critical early weeks of membership. Critical visit counts for each week, weeks 6 through 52, are included in Figure 4(d).
Figure 1: Estimating critical number of visits. We compared the visit distributions of the members who survived less than or equal to 6 weeks and survived more than 6 weeks (a) by looking at their cumulative distributions (b) to determine the critical visit threshold that best separates the two groups (c).
2.5 Causal Inference Framework
The fundamental objective of our investigation is to understand the causal effects of diverse gym-related interventions on the establishment of consistent exercise habits. While traditional correlation-based methodologies provide valuable insights, they remain inherently limited in their capacity to distinguish genuine causal relationships from mere statistical associations. Initially, we employed a supervised learning approach, constructing a logistic regression model to predict six-week survival (defined as whether a member attained the critical visit threshold) as a function of early interventions and relevant covariates (including age, gender, BMI, behavioral cluster, membership details, and self-reported form/experience/frequency metrics). This preliminary model revealed statistically significant positive associations; increased group lesson attendance, more frequent personal trainer sessions, and greater club variability were all correlated with enhanced odds of survival. However, this approach could not effectively disentangle whether these relationships were causally meaningful or merely reflections of confounding variables (for instance, intrinsically motivated individuals might simultaneously select more interventions and demonstrate greater persistence). Furthermore, standard logistic regression treats all variables with symmetrical importance and lacks an explicit mechanism to encode structural assumptions regarding which factors might confound specific treatment–outcome relationships. While valuable for predictive purposes, such supervised methods fundamentally lack the formal identification and rigorous adjustment strategies essential for robust causal interpretation.
To address these methodological limitations, we utilized a causal inference framework based on clear graphical assumptions and techniques for adjusting propensity scores. This approach allows us to estimate the effects of interventions under counterfactual conditions. It effectively measures outcomes as if participants had been randomly assigned to different treatment levels, thereby isolating the true causal impact from misleading correlational artifacts.
We defined our treatment variables $T\in\mathcal{T}$ as systematically binned levels of five distinct gym interventions measured over the initial six-week period: group lessons, personal-trainer sessions, variety of lesson types, diversity of club visits, and invitation credit usage. Additionally, we incorporated three self-reported variables (form level, experience level, and estimated visit frequency) as supplementary treatment indicators. The binary outcome variable $Y\in\{0,1\}$ indicates whether a member successfully achieved the six-week critical-visit milestone. Potential confounding variables $\mathbf{X}\in\mathcal{X}$ encompass demographic factors (age, gender, BMI), membership characteristics (start date, primary club affiliation, membership category), and initial behavioral cluster assignment. These covariates were selected based on their theoretical potential to simultaneously influence both the probability of receiving specific treatments and the likelihood of establishing consistent attendance patterns.
We formalized our causal assumptions through a directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $\mathcal{V}$ represents nodes and $\mathcal{E}$ directed edges. In this graph, edges connect confounders $X$ to both treatment $T$ and outcome $Y$, with additional edges directly linking $T$ to $Y$. Under this causal structure, the average treatment effect (ATE) of setting $T=t$ is identified through the backdoor adjustment formula:
$\mathbb{E}[Y\mid\mathrm{do}(T=t)]$ $=\sum_{x\in\mathcal{X}}\mathbb{E}[Y\mid T=t,\mathbf{X}=x]\,P(\mathbf{X}=x)$ (3)
$=\mathbb{E}_{\mathbf{X}}[\mathbb{E}[Y\mid T=t,\mathbf{X}]]$ (4)
where $\mathrm{do}(T=t)$ represents Pearl’s do-calculus operator indicating intervention rather than passive observation. We verified that conditioning on $\mathbf{X}$ effectively blocks all non-causal backdoor paths between $T$ and $Y$, thereby satisfying the backdoor criterion for causal identification.
To implement the backdoor adjustment efficiently, we estimate propensity scores $e(\mathbf{x})=P(T=1\mid\mathbf{X}=\mathbf{x})$ via logistic regression and execute one-to-one nearest-neighbor matching without replacement, employing the following estimator:
$\hat{e}(\mathbf{x})$ $=\frac{\exp(\mathbf{x}^{T}\boldsymbol{\beta})}{1+\exp(\mathbf{x}^{T}\boldsymbol{\beta})}$ (5)
where $\boldsymbol{\beta}$ represents the vector of regression coefficients. This matching procedure effectively balances the joint distribution of $\mathbf{X}$ between treated and control groups, as confirmed by standardized mean differences consistently below the conventional threshold of 0.1 across all covariates. The resulting matched sample approximates the conditions of a randomized experiment, enabling unbiased estimation of the ATE through:
$\widehat{\mathrm{ATE}}$ $=\frac{1}{N}\sum_{i=1}^{N}\bigl(Y_{i}^{(T=1)}-Y_{i}^{(T=0)}\bigr)$ (6)
$=\frac{1}{|\mathcal{M}|}\sum_{i\in\mathcal{M}}\bigl(Y_{i}-Y_{m(i)}\bigr)$ (7)
where $Y_{i}^{(t)}$ denotes the potential outcome for unit $i$ under treatment level $t$, $\mathcal{M}$ represents the set of treated units that were successfully matched, and $m(i)$ denotes the matched control unit for treated unit $i$.
To establish well-defined treatment conditions, continuous and count-based intervention variables are systematically categorized into four distinct bins: None (no intervention), Low (intervention level $\leq$ 33rd percentile), Moderate (intervention level between 33rd and 66th percentiles), and High (intervention level >> 66th percentile). Self-reported ordinal variables with $n$ levels are binarized using n1n-1 threshold indicators (e.g., ”form level $\geq 1$” versus ”0,” followed by ”form level =2=2” versus $\leq 1$”), thereby preserving the inherent ordering while ensuring adequately balanced sample sizes across treatment conditions.
We used the DoWhy library (Sharma and Kiciman, 2020) to implement our causal inference pipeline through a systematic workflow:
  1. 1.
    Specify the dataset triplet $(Y,T,\mathbf{X})$ and formalize the causal structure via DAG notation.
  2. 2.
    Automatically verify the applicability of backdoor adjustment through d-separation tests.
  3. 3.
    Estimate propensity models and execute matching procedures with appropriate diagnostics.
  4. 4.
    Calculate point estimates of ATEs and construct nonparametric bootstrap confidence intervals.
To ensure the reliability of our causal estimates against potential methodological artifacts or unobserved confounding, we conducted sensitivity analysis as well. We introduce randomly generated variables as pseudo-treatments, confirming that no significant causal effects are detected (yielding $p$-values $\approx 1.0$ as expected under the null hypothesis).
To capture potential heterogeneity in treatment effects across behavioral patterns, we also independently replicated the propensity score matching (PSM) procedure within each of the five NMF-derived behavioral clusters. In these stratified models, we exclude the cluster label from the confounder set $\mathbf{X}$ to prevent collinearity issues, thereby enabling unbiased estimation of cluster-specific treatment effects through:
$\widehat{\mathrm{ATE}}_{c}$ $=\frac{1}{|\mathcal{M}_{c}|}\sum_{i\in\mathcal{M}_{c}}\bigl(Y_{i}-Y_{m_{c}(i)}\bigr)$ (8)
where $\mathcal{M}_{c}$ represents the matched treated units within cluster $c$, and $m_{c}(i)$ denotes the matched control unit for treated unit $i$ within that cluster.
Table 1: Covariates used for confounding adjustment in PSM for each intervention.
Demographic Membership Experience
Intervention Age Gender BMI Start Club Category Form Cluster
Group Lessons
Personal Trainer Sessions
Invitation Credits
Different Club Visits
Different Group Lessons
Self-Reported Variables
Form Level
Experience Level
Estimated Visit Frequency
By integrating preliminary associative modeling with a specified causal inference framework; we derive methodologically sound and reproducible estimates of how early gym interventions causally influence the establishment and maintenance of consistent exercise habits.